20 men finished 1/4 of the house in 7 days.but they have to finish it in 5 more days. how many men shall be added to the crew so that they can finish the house?
Answer by stanbon(75887)

20 men finished 1/4 of the house in 7 days.but they have to finish it in 5 more days. how many men shall be added to the crew so that they can finish the house?
----------------------------
# of men and # of houses completed are directly related
# of men and # of days required are inversely related
--------------
General Equation:
men = k[h/d]
Need to find "k"
---------------
Your problem:
20 = k[(1/4)/7]
k = 28*20 = 560
-----------------
Your Equation:
m = 560[h/d]
Substitute your given values of h= 3/4 of a house and d = 5 days
to get:
m = 560[(3/4)/5]
m = 140*3/5
m = 28*3
m = 84 (# of men to complete 3/4 th of the house in 5 days)
==============
Cheers,
Stan H.